laurie left home and ran to the lake at 10 mi/h. She ran back home at 8 mi/h. if the entire trip took 27 min, how far did she run in all
Answer by stanbon(75887)

laurie left home and ran to the lake at 10 mi/h. She ran back home at 8 mi/h. if the entire trip took 27 min, how far did she run in all
------
To the Lake DATA:
rate = 10 mph ; distance = x miles ; time = x/10 hrs
----
Back to home DATA:
rate = 8 mph ; distance = x miles; time = x/8 hrs
-------
Equation:
time + time = (28/60) hrs
----
x/10 + x/8 = 7/15
-----
15*8x + 15*10x = 7*8*10
-----
120x + 150x = 560
----
270x = 560
------
x = 2.07 miles
-----------
Cheers,
Stan H.
---------